Linux AX25-HOWTO, Amateur Radio.

Terry Dawson, VK2KTJ, terry@perf.no.itg.telecom.com.au v1.2, 26 November 1996 The Linux Operating System is perhaps the only operating system in the world that can boast native and standard support for the AX.25 packet radio protocol utilised by Amateur Radio Operators worldwide. This document aims to describe how to install and configure this support.

Contents

1 Introduction. 4 1.1 Changes from the previous version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2 Other related documentation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Where to obtain new versions of this document. 5

3 Linux and the AX.25 and NetRom Packet Radio Protocols. 5

4 The AX.25/NetRom software components. 5

4.1 Finding the kernel, tools and utility packages. . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.1.1 The kernel source: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.1.2 The network tools: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.1.3 The AX25 utilities: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

5 Installing the AX.25/NetRom software. 7 5.1 Compiling the kernel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 5.2 The network configuration tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5.2.1 Building the standard net-tools release. . . . . . . . . . . . . . . . . . . . . . . . . . . 8 5.3 The AX.25 user programs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

6 Configuring a RAW AX.25 connected mode interface. 10 6.1 Creating the /etc/ax25/axports file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 6.2 Creating the AX.25 device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6.2.1 Setting the callsign for 8530 SCC cards. . . . . . . . . . . . . . . . . . . . . . . . . . . 11 6.2.2 Testing the interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6.3 Configuring for Dual Port TNC’s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 7 Configuring an AX.25 interface for tcp/ip. 13

8 Configuring an AX.25 interface for NetRom. 13 8.1 Configuring /etc/ax25/nrports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 8.2 Configuring /etc/ax25/nrbroadcast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 8.3 Creating the NetRom device . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 8.4 Starting the NetRom daemon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

9 Configuring IP on a NetRom interface. 16

10 Configuring an Ethernet device for AX.25 ala BPQ 16

11 Ottawa PI/PI2 card driver and Gracilis PacketTwin driver. 17

12 Z8530 SCC Kiss Emulation 18 12.1 Obtaining and building the configuration tool package. . . . . . . . . . . . . . . . . . . . . . . 18 12.2 Configuring the driver for your card. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

12.2.1 Configuration of the hardware parameters. . . . . . . . . . . . . . . . . . . . . . . . . 20 12.2.2 Channel Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

12.3 Using the driver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 12.4 The sccstat and sccparam tools. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

13 Configuring an interface for a Baycom modem. 25 13.1 Compiling the kernel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.2 Compiling the setbaycom utility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.3 Creating the /dev device files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.4 Loading the module. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 13.5 A small trap to be aware of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 13.6 Some example configurations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 13.7 Configuring multiple devices. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 13.8 Actually using the devices. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

14 Configuring ax25d to accept incoming AX.25/Netrom connections. 29 14.1 Creating the /etc/ax25/ax25d.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 14.2 A simple example ax25d.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 14.3 Starting ax25d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 14.4 Configuring the node software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

14.4.1 Creating the /usr/local/node/node.conf file. . . . . . . . . . . . . . . . . . . . . . . 33 14.4.2 Creating the /usr/local/node/node.perms file. . . . . . . . . . . . . . . . . . . . . . 34 14.4.3 Configuring node to run from ax25d . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 14.4.4 Configuring node to run from inetd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

14.5 Configuring axspawn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

14.5.1 Creating the /etc/ax25/axspawn.conf file. . . . . . . . . . . . . . . . . . . . . . . . . 37 14.6 Configuring the pms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

14.6.1 Customising pms.c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 14.6.2 Create the /usr/local/pms/motd file. . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 14.6.3 Create the /usr/local/pms/info file. . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 14.6.4 Associate AX.25 callsigns with system users. . . . . . . . . . . . . . . . . . . . . . . . 39 14.6.5 Add the PMS to the /etc/ax25/ax25d.conf file. . . . . . . . . . . . . . . . . . . . . . 40 14.6.6 Test the PMS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

15 HOWTO link NOS and the Linux kernel networking software 40 15.1 Linking NOS and Linux using a pipe device . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

16 The /proc/ file system entries. 42

17 Walking on the wilder side .. the developmental software 42 17.1 Enhanced versions of just about everything. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 17.2 ROSE Packet Layer Support . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 17.3 Soundcard DSP based modem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

18 Some sample configurations. 43 18.1 Small Ethernet LAN with Linux as a router to Radio LAN . . . . . . . . . . . . . . . . . . . 43 18.2 IPIP encapsulated gateway configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

19 Discussion relating to Amateur Radio and Linux. 50

20 Copyright. 50

1 Introduction.

This document was originally an appendix to the HAM-HOWTO, but grew too large to be reasonably managed in that fashion. This document describes how to install and configure the native AX.25 and NetRom support for Linux. A few typical configurations are described that could be used as models to work from.

1.1 Changes from the previous version

Additions:
        Added 'node' configuration.
        Added 'mkiss' configuration.
        Added warning about -assoc 'root' - thanks Jonathon.
        The files in /proc.
        The alpha software.
Corrections/Updates:
        Corrected Baycom MAJOR - thanks John.
        Corrected ax25d mode specification.
        Updated ax25-utils.
        Updated z8530 SCC.
        Fixed munge sample output - thanks Ron.
        Fixed BPQPARAMS - thanks Petteri.
        Updated and added the setbaycom utility.

1.2 Other related documentation.

There is a lot of related documentation. There are many documents that relate to Linux networking in more general ways and I strongly recommend you also read these as they will assist you in your efforts and provide you with stronger insight into other possible configurations.

They are:

The HAM-HOWTO <http://sunsite.unc.edu/mdw/HOWTO/HAM-HOWTO.html>
the Net-2-HOWTO <http://sunsite.unc.edu/mdw/HOWTO/NET-2-HOWTO.html>
the Ethernet-HOWTO <http://sunsite.unc.edu/mdw/HOWTO/Ethernet-HOWTO.html>
the IPX-HOWTO <http://sunsite.unc.edu/mdw/HOWTO/IPX-HOWTO.html>

and

the Firewall-HOWTO <http://sunsite.unc.edu/mdw/HOWTO/Firewall-HOWTO.html>

2 Where to obtain new versions of this document.

The best place to obtain the latest version of this document is from a Linux Documentation Project archive. The Linux Documentation Project runs a Web Server and this document appears there as The

AX25-HOWTO <http://sunsite.unc.edu/mdw/HOWTO/AX25-HOWTO.html>. You can always contact me,

but I pass new versions of the document directly to the LDP HOWTO coordinator, so if it isn’t there then chances are I haven’t finished it.

3 Linux and the AX.25 and NetRom Packet Radio Protocols.

The AX.25 protocol offers both connected and connectionless modes of operation, and is used either by itself for point-point links, or to carry other protocols such as TCP/IP and NetRom.

It is similar to X.25 level 2 in structure, with some extensions to make it more useful in the amateur radio environment.

The NetRom protocol is an attempt at a full network protocol and uses AX.25 at its lowest layer as a datalink protocol. It provides a network layer that is an adapted form of AX.25.

Alan Cox developed some early kernel based AX.25 software support for Linux. Jonathon Naylor <g4klx@g4klx.demon.co.uk> has taken up ongoing development of the code, has added NetRom support and is now the developer of the AX.25 related kernel code and utilities. DAMA support was developed by Joerg, DL1BKE, jreuter@lykos.tng.oche.de. Finally Baycom support was added by Thomas Sailer,

<sailer@ife.ee.ethz.ch>.

The Linux code supports KISS based TNC’s (Terminal Node Controllers), the Ottawa PI card, the Gracilis PacketTwin card and other Z8530 SCC based cards with the generic SCC driver and both the Parallel and Serial port Baycom modems.

The User programs contain a simple PMS (Personal Message System), a beacon facility, a line mode connect program, ‘listen’ an example of how to capture all AX.25 frames at raw interface level and programs to configure the NetRom protocol. Included also are an AX.25 server style program to handle and despatch incoming AX.25 connections and a NetRom daemon which does most of the hard work for NetRom support.

4 The AX.25/NetRom software components.

The AX.25 software is comprised of three components, the kernel source, the network configuration tools and the utility programs.

The version 2.0.xx Linux kernels include the AX.25, NetRom, Z8530 SCC, PI card and PacketTwin drivers by default. I recommend you obtain and use the version 2.0.xx kernel source.

4.1 Finding the kernel, tools and utility packages.

4.1.1 The kernel source:

The kernel source can be found in its usual place at: ftp.funet.fi

     /pub/Linux/PEOPLE/Linus/v2.0/

It is stronly recommended that you upgrade to the most recent 2.0 kernel. Jonathon specifically recommends not using a kernel less than 2.0.12 for AX.25 purposes as some minor bugs were fixed in that and later versions.

4.1.2 The network tools:

The latest alpha release of the standard Linux network tools support AX.25 and NetRom and can be found at: ftp.inka.de

     /pub/comp/Linux/networking/net-tools/net-tools-1.32-alpha.tar.gz

or: ftp.linux.org.uk

     /pub/linux/Networking/PROGRAMS/NetTools/net-tools-1.32-alpha.tar.gz

The latest ipfwadm package can be found at: ftp.xos.nl

     /pub/linux/ipfwadm/

4.1.3 The AX25 utilities:

The AX.25 utility programs can be found at their home site:

Jonathon Naylor’s Pages <http://www.cs.nott.ac.uk/ jsn/>

or by ftp from: ftp.cs.nott.ac.uk

     /jsn/ax25-utils-2.0.12a.tar.gz

or: sunsite.unc.edu

     /pub/Linux/apps/ham/ax25-utils-2.0.12a.tar.gz

5 Installing the AX.25/NetRom software.

To successfully install AX.25 support on your linux system you must configure and install an appropriate kernel and then install the AX.25 utilities.

5.1 Compiling the kernel.

If you are already familiar with the process of compiling the Linux Kernel then you can skip this section, just be sure to select the appropriate options when compiling the kernel. If you are not, then read on.

The normal place for the kernel source to be unpacked to is the /usr/src directory into a subdirectory called linux. To do this you should be logged in as root and execute a series of commands similar to the following:

     # mv linux linux.old
     # cd /usr/src
     # gzip -dc linux-2.0.25.tar.gz | tar xvofp -
     # cd linux

After you have unpacked the kernel source into place you need to run the configuration script and choose the options that suit your hardware configuration and the options that you wish built into your kernel. You do this by using the command:

     # make config

You might also try:

     # make menuconfig

if you prefer a full screen menu based method. I’m going to describe the original method, but you use whichever you are most comfortable with.

In either case you will be offered a range of options at which you must answer ‘Y’ or ‘N’. (Note you may also answer ‘M’ if you are using modules. For the sake of simplicity I will assume you are not, please make appropriate modifications if you are).

The options most relevant to an AX.25 configuration are:

Prompt for development and/or incomplete code/drivers [Y/n]         Y
Networking support (CONFIG_NET) [Y/n/?]                             Y
Loopback device support (CONFIG_BLK_DEV_LOOP) [N/y/m/?]             Y
Network firewalls (CONFIG_FIREWALL) [N/y/?]
TCP/IP networking (CONFIG_INET) [Y/n/?]                             Y
IP: forwarding/gatewaying (CONFIG_IP_FORWARD) [N/y/?]
IP: firewalling (CONFIG_IP_FIREWALL) [N/y/?] (NEW)
IP: tunneling (CONFIG_NET_IPIP) [N/y/m/?] (NEW)

Amateur Radio AX.25 Level 2 (CONFIG_AX25) [N/y/?]                   Y
AX.25 over Ethernet (CONFIG_BPQETHER) [N/y/?] (NEW)
Amateur Radio NET/ROM (CONFIG_NETROM) [N/y/?] (NEW)
Kernel/User network link driver(ALPHA) (CONFIG_NETLINK) [N/y/?]
Network device support (CONFIG_NETDEVICES) [Y/n/?]                  Y
SLIP (serial line) support (CONFIG_SLIP) [N/y/m/?]                  Y
Radio network interfaces (CONFIG_NET_RADIO) [N/y/?]
BAYCOM ser12 and par96 kiss emulation driver for AX.25 (CONFIG_BAYCOM) [N/y/m/?]
Z8530 SCC kiss emulation driver for AX.25 (CONFIG_SCC) [N/y/m/?]
Other ISA cards (CONFIG_NET_ISA) [N/y/?]
Ottawa PI and PI/2 support (CONFIG_PI) [N/y/?] (NEW)
Gracilis PackeTwin support (CONFIG_PT) [N/y/?] (NEW)
Standard/generic serial support (CONFIG_SERIAL) [Y/n/?]

The options I have flagged with a ‘Y’ are those that you must must answer ‘Y’ to. The rest are dependent on what hardware you have and what options you want to include. Some of these options are described in more detail later on, so if you don’t know what you want yet, then read ahead and come back to this step later.

After you have completed the kernel configuration you should be able to cleanly compile your new kernel:

     # make dep
     # make clean
     # make zImage

Make sure you move your arch/i386/boot/zImage file wherever you want it and then run lilo to ensure that you actually boot from it.

5.2 The network configuration tools.

After you have compiled the kernel you should compile the new network configuration tools. These tools allow you to modify the configuration of network devices and to add routes to the routing table.

The new alpha release of the standard net-tools package includes support for AX.25 and NetRom support. I’ve tested this and it seems to work well for me.

5.2.1 Building the standard net-tools release.

Don’t forget to read the Release file and follow any instructions there. The steps I used to compile the tools were:

     # cd /usr/src
     # tar xvfz net-tools-1.32-alpha.tar.gz
     # cd net-tools-1.32-alpha
     # make config

At this stage you will be presented with a series of configuration questions, similar to the kernel configuration questions. Be sure to include support for all of the protocols and network devices types that you intend to use. If you do not know how to answer a particular question then answer ‘Y’.

The tools compiled cleanly against a 2.0.0 kernel without error.

When the compilation is complete, you should need only use the:

     # make install

command to install the programs in their proper place.

If you wish to use the IP firewall facilities then you will need the latest firewall administration tool ipfwadm. This tool replaces the older ipfw tool which will not work with new kernels.

I compiled the ipfwadm utility with the following commands:

     # cd /usr/src
     # tar xvfz ipfwadm-2.0beta2.tar.gz
     # cd ipfwadm-2.0beta2
     # make install
     # cp ipfwadm.8 /usr/man/man8
     # cp ipfw.4 /usr/man/man4

5.3 The AX.25 user programs.

After you have successfully compiled and booted your new kernel, you need to compile the user programs. To compile and install the user programs you should use a series of commands similar to the following:

     # cd /usr/src
     # gzip -dc ax25-utils-2.0.12a.tar.gz | tar xvvof -
     # cd ax25-utils-2.0.12a
     # ./configure
     # make
     # make install

The files will be installed under the /usr directory by default in subdirectories: bin, sbin, etc and man.

If you get messages something like:

gcc -Wall -Wstrict-prototypes -O2 -I../lib -c call.c
call.c: In function `statline':
call.c:268: warning: implicit declaration of function `attron'
call.c:268: `A_REVERSE' undeclared (first use this function)
call.c:268: (Each undeclared identifier is reported only once
call.c:268: for each function it appears in.)

then you should read the README.ncurses file in the top level directory. You can do as Jonathon suggests in the INSTALL file, or you can try replacing:

#include <curses.h>

with:

#include <ncurses/curses.h>

in the:

call/call.c
call/menu.c

files.

6 Configuring a RAW AX.25 connected mode interface.

The first stage to configuration of an AX.25 interface is to configure it to work as a ‘vanilla’ AX.25 interface with no TCP/IP. The following configuration will get you to the point of being able to make AX.25 calls from your Linux machine to other AX.25 nodes.

The AX.25 software has been designed to work with a TNC in kiss mode or with other cards such as the Ottawa PI2 card, PacketTwin and other SCC cards via special drivers that emulate a kiss tnc.

For real KISS TNC’s there are two steps to complete in order to create an AX.25 port ready to use to make outgoing calls. For other types of hardware you need only complete the first step.

6.1 Creating the /etc/ax25/axports file.

The AX.25 ports have a configuration file that is read by many programs that want to find information about an AX.25 port. This file is called the:

     /etc/ax25/axports

file. The format of the file is as follows:

     portname  callsign  baudrate  paclen  window  description

where:

portname is a text name that you will refer to the port by.

callsign is the AX.25 callsign you want to assign to the port. baudrate

is the speed at which you wish the port to communicate with your TNC.

paclen is the maximum packet length you want to configure the port to use for AX.25 connected mode connections.

window is the AX.25 window (K) parameter. This is the same as the MAXFRAME setting of many tnc’s.

description is a textual description of the port.

In my case, mine looks like:

     radio    VK2KTJ-0        4800    256     2       144.800 MHz

6.2 Creating the AX.25 device

If you are using an SCC card like the PI2 or PacketTwin then you do not need to create the network device, as the kernel driver will automatically do this for you. If you are using a KISS TNC then you will need to create the AX.25 interface as it will not already exist. Creating an AX.25 port is very similar to creating a slip device.

You will need to have the TNC preconfigured and connected to your serial port. You can use a communications program like minicom or seyon to configure the TNC into kiss mode if you wish.

In it simplest form you can use the axattach program as follows:

     # /usr/sbin/axattach /dev/ttyS0 radio

would configure your /dev/ttyS0 serial device to be a kiss interface configured as per the details for the line beginning with the portname ”radio” in the /etc/ax25/axports file.

All this step has done is to actually activate the device in the kernel, you need to run other programs before you an actually make use of the port.

6.2.1 Setting the callsign for 8530 SCC cards.

If you are using a PI or PacketTwin card then you should use the axparms -setcall command to change the callsign of the appropriate port to that which you intend to use. Refer to the PI/PacketTwin section for the names of the network devices to use.

6.2.2 Testing the interface.

You now should be able to make outgoing AX.25 connections. To test AX.25 connected mode you could use the call program as demonstrated:

     /usr/bin/call radio VK2DAY via VK2SUT

Note: you must tell call which AX.25 port you wish to make the call on, as the same AX.25 node might be reachable on any of the ports you have configured.

The call program is a linemode terminal program for making AX.25 calls. It recognises lines that start with ‘~’ as command lines. The ‘~.’ command will close the connection.

Please refer to the man page in /usr/man and the README file in ax25-utils-2.0.12a distribution for more information.

6.3 Configuring for Dual Port TNC’s

The mkiss utility included in the ax25-utils distribution allows you to make use of both modems on a dual port TNC. Configuration is fairly simple. It works by taking a single serial device connected to a single multiport TNC look like a number of devices each connected to a single port TNC. You do this before you do any of the AX.25 configuration. The devies that you then do the AX.25 configuration on are pseudo-TTY interfaces, (/dev/ttyp*), and not the actual serial device.

Example: if you have a dual port tnc and it is connected to your /dev/ttyS0 serial device at 9600 bps, the command:

     # mkiss -s 9600 /dev/ttyS0 /dev/ptyp0 /dev/ptyp1
     # /usr/sbin/axattach /dev/ttyp0 port1
     # /usr/sbin/axattach /dev/ttyp1 port2

would create two tty devices: /dev/ttyp0 and /dev/ttyp0 that each look like a normal single port TNC. You would then treat /dev/ttyp0 and /dev/ttyp1 just as you would a conventional serial device with TNC connected. This means you’d then use the axattach command as described above, on each of those, in the example for AX.25 ports called port1 and port2. You shouldn’t use axattach on the actual serial device as the mkiss program uses it.

The mkiss command has a number of optional arguments that you may wish to use. They are summarised as follows:

-c enables the addition of a one byte checksum to each KISS frame. This is not supported by most KISS implementation, it is supported by the G8BPG KISS rom.

-s <speed¿ sets the speed of the serial port.

-h enables hardware handshaking on the serial port, it is off by default. Most KISS implementation do not support this, but some do.

-l enables logging of information to the syslog logfile.

7 Configuring an AX.25 interface for tcp/ip.

It is very simple to configure an AX.25 port to carry tcp/ip. Jonathon has modified the axattach command so that you can include the IP address that you want the interface to be configured with on the command line. For example, to modify the example presented earlier, I could use the command:

     # /usr/sbin/axattach -i 44.136.8.5 -m 512 /dev/ttyS0 radio
     # /sbin/route add -net 44.136.8.0 sl0
     # /sbin/route add default sl0

to create the AX.25 interface with an IP address of 44.136.8.5 and an MTU of 512 bytes. You should still use the ifconfig to configure the other parameters if necessary.

You do not need to do it this way, you can of course just use the ifconfig program to configure the ip address and netmask details for the port and add a route via the port, just as you would for any other tcp/ip interface.

     # /sbin/ifconfig sl0 44.136.8.5
     # /sbin/ifconfig sl0 netmask 255.255.255.0
     # /sbin/ifconfig sl0 broadcast 44.136.8.255
     # /sbin/ifconfig sl0 arp mtu 256 up
     # /sbin/route add -net 44.136.8.0 sl0
     # /sbin/route add default sl0

The commands listed above are typical of the sort of configuration many of you would be familiar with if you have used NOS or any of its derivatives or any other tcp/ip software. Note that the default route might not be required in your configuration if you have some other network device configured.

To test it out, try a ping or a telnet to a local host.

     # ping -i 5 44.136.8.58

Note the use of the ‘-i 5’ arguments to ping to tell it to send pings every 5 seconds instead of its default of 1 second.

8 Configuring an AX.25 interface for NetRom.

To configure NetRom on an AX.25 interface you must configure two files.

8.1 Configuring /etc/ax25/nrports

The first is the /etc/ax25/nrports file. This file describes the NetRom port in much the same way as the /etc/ax25/axports file describes the AX.25 ports. This file is formatted as follows:

     name callsign  alias  paclen   description

Where:

name

is the text name that you wish to refer to the port by. callsign

is the callsign that the NetRom traffic from this port will use. alias

is the NetRom alias this port will have assigned to it. paclen

is the maximum size of NetRom frames transmitted by this port. description

is a free text description of the port.

An example would look something like the following:

     netrom  VK2KTJ-9        LINUX   235     Linux Switch Port

Note that the ports are referred to by name by programs such as call.

8.2 Configuring /etc/ax25/nrbroadcast
The second file is the /etc/ax25/nrbroadcast file.

This file is formatted as follows:

     axport min_obs def_qual worst_qual verbose

Where:

axport is the port name obtained from the /etc/ax25/axports file. If you do not have an entry in /etc/ax25/nrbroadcasts for a port then this means that no NetRom routing will occur and any received NetRom broadcasts will be ignored for that port. min obs

is the minimum obselesence value for the port.

def qual is the default quality for the port.

worst qual is the worst quality value for the port, any routes under this quality will be ignored.

verbose is a flag determining whether full NetRom routing broadcasts will occur from this port or only a routing broadcast advertising the node itself.

An example would look something like the following:

     radio    1       200      10         1

8.3 Creating the NetRom device

When you have the two configuration files completed you must create the NetRom device in much the same way as you did for the AX.25 devices. This time you use the nrattach command:

     # nrattach netrom

This command would start the NetRom device (nr0) named netrom configured with the details specified in

the /etc/ax25/nrports file.

8.4 Starting the NetRom daemon

The NetRom daemon manages the NetRom routing tables and generates the NetRom routing broadcasts. You start NetRom with the command:

     # /usr/sbin/netromd

You should soon see the /proc/net/nr neigh file filling up with information about your NetRom neighbours.

Remember to put the /usr/sbin/netromd command in your rc files so that it is started automatically each time you reboot.

9 Configuring IP on a NetRom interface.

Configuring a NetRom interface for tcp/ip is almost identical to configuring an AX.25 interface for tcp/ip.

Again you can either specify the ip address and mtu on the nrattach command line, or use the ifconfig and route commands, but you need to manually add arp entries for hosts you wish to route to because there is no mechanism available for your machine to learn what NetRom address it should use to reach a particular IP host.

So, to create an nr0 device with an IP address of 44.136.8.5, an mtu of 512 and configured with the details from the /etc/ax25/nrports file for a NetRom port named netrom you would use:

     # /usr/sbin/nrattach -i 44.136.8.5 -m 512 netrom
     # route add 44.136.8.5 nr0

or you could use something like the following commands manually:

     # ifconfig nr0 44.136.8.5 hw netrom VK2KTJ-9
     # route add 44.136.8.5 nr0

Then for each IP host you wish to reach via NetRom you need to set route and arp entries. To reach a destination host with an IP address of 44.136.80.4 at NetRom address BBS:VK3BBS via a NetRom neighbour with callsign VK2SUT-0 you would use commands as follows:

     # route add 44.136.80.4 nr0
     # arp -t netrom -s 44.136.80.4 vk2sut-0
     # nrparms -nodes vk3bbs + BBS 120 6 sl0 vk2sut-0

The ‘120’ and ‘6’ arguments to the nrparms command are the NetRom quality and obsolescence count values for the route.

10 Configuring an Ethernet device for AX.25 ala BPQ

Linux support BPQ Ethernet compatibility. This enables you to run the AX.25 protocol over your Ethernet LAN and to interwork your linux machine with other BPQ machines on the LAN.

Configuration is quite straightforward. You firstly must have configured a standard Ethernet device. This means you will have compiled your kernel to support your Ethernet card, and executed configuration commands similar to the following:

     # ifconfig eth0 44.136.8.5 netmask 255.255.255.248 up
     # ifconfig eth0 broadcast 44.136.8.103
     # route add -net 44.136.8.96 netmask 255.255.255.248 eth0

If you don’t actually want to run tcp/ip on your ethernet device then all you need to do is use the:

     # ifconfig eth0 up

command to activate the interface.

To configure the BPQ support you need to associate an AX.25 callsign with your Ethernet hardware address. You should use the following:

     # axparms -dev eth0 vk2ktj-9

This says that AX.25 callsign vk2ktj-9 is associated with the eth0 device. Note you should never associate a callsign with the root account as this can cause configuration problems in other programs.

BPQ Ethernet normally uses a multicast address. The Linux implementation does not, and instead it uses the normal Ethernet broadcast address. The NET.CFG file for the BPQ ODI driver should therefore look similar to this:

     LINK SUPPORT
             MAX STACKS 1
             MAX BOARDS 1
     LINK DRIVER E2000                    ; or other MLID to suit your card
             INT 10                       ;
             PORT 300                     ; to suit your card
             FRAME ETHERNET_II
             PROTOCOL BPQ 8FF ETHERNET_II ; required for BPQ - can change PID
     BPQPARAMS                            ; optional - only needed if you want
                                          ; to override the default target addr
             ETH_ADDR  FF:FF:FF:FF:FF:FF  ; Target address

11 Ottawa PI/PI2 card driver and Gracilis PacketTwin driver.

The Ottawa PI and the Gracilis PacketTwin are Z8530 SCC based card for IBM PC type machines that are in common usage by Amateur Radio operators worldwide. The PI card supports a high speed half duplex (single DMA channel) port, and a low speed (<9k6bps interrupt driven) half duplex port. The PI2 is a new version of the card that supports an on board radio modem and improved hardware design. The PacketTwin card supports two high speed port and is capable of accepting on board modems.

The PI card driver was written by David Perry, <dp@hydra.carleton.edu> The PacketTwin card

driver was written by Craig Small VK2XLZ, <csmall@triode.apana.org.au>. Both drivers are a

standard feature of the 1.3.xx kernel.

Please refer to the AX.25 section above for details on how to configure the kernel to include the PI card or the PacketTwin card drivers.

Once you have the appropriate driver configured into your kernel then you should reboot and when you do, you should see mention of the PI or PacketTwin card driver in the messages that appear on the screen while the kernel is booting. The drivers test each of the I/O port addresses that the card might be configured for and reports any that it finds. You should then look at the /proc/net/dev file and you should see reference to devices called pi0a and pi0b or pt0a and pt0b for the PI and PacketTwin cards respectively.

If the devices don’t appear then recheck that you have configured and compiled your kernel correctly and that you are in fact actually running your new kernel. If you still don’t then this suggests that your PI or PacketTwin card was not detected by the kernel while it booted and may indicate that you have some sort of hardware conflicting with your PI/PT card that prevented it being detected.

If all of the above went as planned then you will need to configure your PI/PT card with an AX.25 address and for IP. The configuration of the PI/PT cards is virtually identical to that of any other IP interface. Something like the following should work ok for you:

     # /usr/sbin/axparms -setcall pi0a VK2KTJ-2
     # /sbin/ifconfig pi0a 44.136.8.5
     # /sbin/ifconfig pi0a netmask 255.255.255.0
     # /sbin/ifconfig pi0a broadcast 44.136.8.255
     # /sbin/ifconfig pi0a arp mtu 256 up
     # /sbin/route add -net 44.136.8.0 pi0a
     # /sbin/route add default pi0a

Note that pi0a refers to the ‘a’ port on the first PI card found, and that pi0b would therefore refer to the ‘b’ port on the first PI card found. pt0a would therefore be the first port on the first PacketTwin card detected.

Note also the use of the axparms -setcall command to set the AX.25 callsign of the port.

As usual, when this has been done you can test the interface with the ping or telnet command to ensure it is working.

12 Z8530 SCC Kiss Emulation

Joerg Reuter, DL1BKE, jreuter@lykos.tng.oche.de has developed generic support for Z8530 SCC based cards. His driver is configurable to support a range of different types of cards and present an interface that looks like a KISS TNC so you can treat it as though it were a KISS TNC.

12.1 Obtaining and building the configuration tool package.

While the kernel driver is included in the standard kernel distribution, Joerg distributes more recent versions of his driver with the suite of configuration tools that you will need to obtain as well.

You can obtain the configuration tools package from: db0bm.automation.fh-aachen.de

     /incoming/dl1bke/

or:

insl1.etec.uni-karlsruhe.de

     /pub/hamradio/linux/z8530/

or:

ftp.ucsd.edu

     /hamradio/packet/tcpip/linux
     /hamradio/packet/tcpip/incoming/

You will find multiple versions, choose the one that best suits the kernel you intend to use:

z8530drv-2.4.dl1bke.tar.gz 2.0.* z8530drv-utils-3.0.tar.gz 2.1.6 or greater

The following command were what I used to compile and install the package for kernel version 2.0.25:

     # cd /usr/src
     # gzip -dc z8530drv-2.4.dl1bke.tar.gz | tar xvpofz -
     # cd z8530drv
     # make clean
     # make dep
     # make module         # If you want to build the driver as a module
     # make for_kernel     # If you want the driver to built into your kernel
     # make install

After the above is complete you should have three new programs installed in your /sbin directory: gencfg, sccinit and sccstat. It is these programs that you will use to configure the driver for your card.

You will also have a group of new special device files created in your /dev called scc0 .. scc7. These will be used later and will be the ‘KISS’ devices you will end up using.

If you chose to ’make for kernel’ then you will need to recompile your kernel. To ensure that you include support for the z8530 driver you must be sure to answer ‘Y’ to: ‘Z8530 SCC kiss emulation driver for AX.25’ when asked during a kernel ‘make config’.

If you chose to ’make module’ then the new scc.o will have been installed in the appropriate /lib/modules directory and you do not need to recompile your kernel. Remember to use the insmod command to load the module before your try and configure it.

12.2 Configuring the driver for your card.

The z8530 SCC driver has been designed to be as flexible as possible so as to support as many different types of cards as possible. With this flexibility has come some cost in configuration.

There is more comprehensive documentation in the package and you should read this if you have any problems. You should particularly look at doc/scc eng.doc or doc/scc ger.doc for more detailed information. I’ve paraphrased the important details, but as a result there is a lot of lower level detail that I have not included.

The main configuration file is read by the sccinit program and is called /etc/z8530drv.rc. This file is broken into two main stages: Configuration of the hardware parameters and channel configuration. After you have configured this file you need only add:

     # sccinit

into the rc file that configures your network and the driver will be initialised according to the contents of the configuration file. You must do this before you attempt to use the driver.

12.2.1 Configuration of the hardware parameters.

The first section is broken into stanzas, each stanza representing an 8530 chip. Each stanza is a list of keywords with arguments. You may specify up to four SCC chips in this file by default. The #define MAXSCC 4 in scc.c can be increased if you require support for more.

The allowable keywords and arguments are:

chip the chip keyword is used to separate stanzas. It will take anything as an argument. The arguments are not used.

data a this keyword is used to specify the address of the data port for the z8530 channel ‘A’. The argument is a hexadecimal number e.g. 0x300

ctrl a this keyword is used to specify the address of the control port for the z8530 channel ‘A’. The arguments is a hexadecimal number e.g. 0x304

data b this keyword is used to specify the address of the data port for the z8530 channel ‘B’. The argument is a hexadecimal number e.g. 0x301

ctrl b this keyword is used to specify the address of the control port for the z8530 channel ‘B’. The arguments is a hexadecimal number e.g. 0x305

irq this keyword is used to specify the IRQ used by the 8530 SCC described in this stanza. The argument is an integer e.g. 5 pclock

this keyword is used to specify the frequency of the clock at the PCLK pin of the 8530. The argument is an integer frequency in Hz which defaults to 4915200 if the keyword is not supplied.

board the type of board supporting this 8530 SCC. The argument is a character string. The allowed values are: PA0HZP

the PA0HZP SCC Card EAGLE

the Eagle card PC100

the DRSI PC100 SCC card PRIMUS

the PRIMUS-PC (DG9BL) card BAYCOM

BayCom (U)SCC card

escc this keyword is optional and is used to enable support for the Extended SCC chips (ESCC) such as the 8580, 85180, or the 85280. The argument is a character string with allowed values of ‘yes’ or ‘no’. The default is ‘no’.

vector this keyword is optional and specifies the address of the vector latch (also known as ”intack port”) for PA0HZP cards. There can be only one vector latch for all chips. The default is 0.

special this keyword is optional and specifies the address of the special function register on several cards. The default is 0.

option this keyword is optional and defaults to 0.

Some example configurations for the more popular cards are as follows:

BayCom USCC

     chip    1
     data_a  0x300
     ctrl_a  0x304
     data_b  0x301
     ctrl_b  0x305

12. Z8530 SCC Kiss Emulation 22

irq     5
board   BAYCOM
#
# SCC chip 2
#
chip    2
data_a  0x302
ctrl_a  0x306
data_b  0x303
ctrl_b  0x307
board   BAYCOM

PA0HZP SCC card

chip 1
data_a 0x153
data_b 0x151
ctrl_a 0x152
ctrl_b 0x150
irq 9
pclock 4915200
board PA0HZP
vector 0x168
escc no
#
#
#
chip 2
data_a 0x157
data_b 0x155
ctrl_a 0x156
ctrl_b 0x154
irq 9
pclock 4915200
board PA0HZP
vector 0x168
escc no

DRSI SCC card

chip 1
data_a 0x303
data_b 0x301
ctrl_a 0x302
ctrl_b 0x300
irq 7
pclock 4915200
board DRSI
escc no

If you already have a working configuration for your card under NOS, then you can use the gencfg command to convert the PE1CHL NOS driver commands into a form suitable for use in the z8530 driver configuration file.

To use gencfg you simply invoke it with the same parameters as you used for the PE1CHL driver in NET/NOS. For example:

     # gencfg 2 0x150 4 2 0 1 0x168 9 4915200

will generate a skeleton configuration for the OptoSCC card.

12.2.2 Channel Configuration

The Channel Configuration section is where you specify all of the other parameters associated with the port you are configuring. Again this section is broken into stanzas. One stanza represents one logical port, and therefore there would be two of these for each one of the hardware parameters stanzas as each 8530 SCC supports two ports.

These keywords and arguments are also written to the /etc/z8530drv.rc file and must appear after the hardware parameters section.

Sequence is very important in this section, but if you stick with the suggested sequence it should work ok. The keywords and arguments are:

device this keyword must be the first line of a port definition and specifies the name of the special device file that the rest of the configuration applies to. e.g. /dev/scc0

speed this keyword specifies the speed in bits per second of the interface. The argument is an integer: e.g.

     1200

clock this keyword specifies where the clock for the data will be sourced. Allowable values are:

dpll

normal halfduplex operation external

MODEM supplies its own Rx/Tx clock divider

use fullduplex divider if installed.

mode this keyword specifies the data coding to be used. Allowable arguments are: nrzi or nrz rxbuffers

this keyword specifies the number of receive buffers to allocate memory for. The argument is an integer, e.g. 8.

txbuffers this keyword specifies the number of transmit buffers to allocate memory for. The argument is an integer, e.g. 8.

bufsize this keyword specifies the size of the receive and transmit buffers. The arguments is in bytes and represents the total length of the frame, so it must also take into account the AX.25 headers and not just the length of the data field. This keyword is optional and default to 384

txdelay

the KISS transmit delay value, the argument is an integer in mS. persist

the KISS persist value, the argument is an integer. slot

the KISS slot time value, the argument is an integer in mS. tail

the KISS transmit tail value, the argument is an integer in mS. fulldup

the KISS full duplex flag, the argument is an integer. 1==Full Duplex, 0==Half Duplex. wait

the KISS wait value, the argument is an integer in mS. min

the KISS min value, the argument is an integer in S. maxkey

the KISS maximum keyup time, the argument is an integer in S. idle

the KISS idle timer value, the argument is an integer in S. maxdef

the KISS maxdef value, the argument is an integer. group

the KISS group value, the argument is an integer. txoff

the KISS txoff value, the argument is an integer in mS.

softdcd the KISS softdcd value, the argument is an integer.

slip the KISS slip flag, the argument is an integer.

12.3 Using the driver.

To use the driver you simply treat the /dev/scc* devices just as you would a serial tty device with a KISS TNC connected to it. For example, to configure Linux Kernel networking to use your SCC card you could use something like:

     # axattach -s 4800 /dev/scc0 VK2KTJ

You can also use NOS to attach to it in precisely the same way. From JNOS for example you would use something like:

     attach asy scc0 0 ax25 scc0 256 256 4800

12.4 The sccstat and sccparam tools.

To assist in the diagnosis of problems you can use the sccstat program to display the current configuration of an SCC device. To use it try:

     # sccstat /dev/scc0

you will displayed a very large amount of information relating to the configuration and health of the

/dev/scc0 SCC port.

The sccparam command allows you to change or modify a configuration after you have booted. Its syntax is very similar to the NOS param command, so to set the txtail setting of a device to 100mS you would use:

     # sccparam /dev/scc0 txtail 0x8

13 Configuring an interface for a Baycom modem.

Thomas Sailer, <sailer@ife.ee.ethz.ch>, despite the popularly held belief that it would not work very well, has developed Linux support for Baycom modems. His driver supports the Ser12 serial port, Par96 and the enhanced Par97 parallel port modems. Further information about the modems themselves may be obtained from the Baycom Web site <http://www.baycom.de>.

When the driver is running it follows the standard of looking like a KISS interface, so you treat it in much the same way as you would the PI card, PacketTwin or SCC driver ports.

13.1 Compiling the kernel

The driver is only available as a module, so you must have compiled module support into your kernel, and also selected the appopriate options to support the Baycom driver during the make config stage when compiling your kernel.

The relevant options that you must answer ‘Y’ for are:

     Prompt for development and/or incomplete code/drivers [Y/n]
     Radio network interfaces (CONFIG_NET_RADIO) [N/y/?]
     BAYCOM ser12 and par96 kiss emulation driver for AX.25 (CONFIG_BAYCOM) [N/y/m/?]

13.2 Compiling the setbaycom utility

A configuration utility is available available from Thomas Sailer’s home page

<http://www.ife.ee.ethz.ch/ sailer/ham/baycom.tar.gz>. The setbaycom utility allows you to

configure multiple Baycom devices. If you intend to do this then you should obtain it and install it.

To compile and install, do the following:

     # cd /usr/src
     # tar xvfz baycom.tar.gz
     # cd baycom
     # make
     # make install

13.3 Creating the /dev device files.

The Baycom driver uses special device files in the /dev directory. You need to create these manually. The following commands should work ok for you:

     # mknod /dev/bc0 c 51 0
     # mknod /dev/bc1 c 51 1
     # mknod /dev/bc2 c 51 2
     # mknod /dev/bc3 c 51 3

You will use these later. If you installed the setbaycom utility then this will have been done for you.

13.4 Loading the module.

When the kernel compilation is complete and you have installed the module baycom.o wherever you normally keep your modules (conventionally they are stored in /lib/modules/2.0.0/), you must start the module. The module takes configuration paramters from the insmod command line, so you must start it manually from an rc file or if you use the kerneld program you must configure it with the relevant paramters so that it supplies them when it loads the module.

If you use kerneld you will need to add the following line to your /etc/conf.modules file:

alias char-major-51 baycom

There are four main items to specify when loading the module, they are:

modem

the baycom modem type you are configuring it for. 1. Ser12 type modem.

2. Par96 or Par97 type modem. iobase

the i/o address of the serial or parallel port you will use. 0x3f8

COM1: 0x2f8

COM2: 0x378

LPT1: 0x278

LPT2:

irq the Interrupt Request (IRQ) line your port will use. 4

COM1: 3

COM2: 7

LPT1: 5

LPT2:

options some baycom specific options. 0

use hardware generated DCD signal. 1

use software DCD signal.

13.5 A small trap to be aware of.

Because the Baycom driver uses the serial and parallel ports of your machines, and because it is a low level driver for these, it competes with the existing serial and parallel drivers in the kernel. There are some ways of working around this problem. The first way is to make sure that you load the Baycom module first so that it takes the hardware it needs before the serial or parallel device driver modules load and search for their own. With the serial driver you have the luxury of being able to disable it for a single port using the setserial command, this will be demonstrated in the examples. If you are using the parallel port modems then beware of the lp.o Line Printer and plip.o Parallel Line IP modules.

13.6 Some example configurations.

These examples are for the two most common configurations.

Disable the serial driver for COM1: then configure the Baycom driver for a Ser12 serial port modem on COM1: with the software DCD option enabled:

     # setserial /dev/ttyS0 uart none
     # insmod baycom modem=1 iobase=0x3f8 irq=4 options=1

Par96 parallel port type modem on LPT1: using hardware DCD detection:

     # insmod baycom modem=2 iobase=0x378 irq=7 options=0

The first device created is /dev/bc0, the second is /dev/bc1 etc.

13.7 Configuring multiple devices.

Using the insmod program and its command line arguments you can only configure one Baycom modem. The setbaycom utility allows you to configure multiple Baycom devices.

The command line arguments needed are identical to those you’d supply to the module if you were configuring just a single device. For example, to configure both of the devices presented above you would use the commands:

     # setbaycom -i /dev/bc0 -p ser12 0x3f8 4 1
     # setbaycom -i /dev/bc1 -p par96 0x378 7 0

13.8 Actually using the devices.

This device driver when loaded looks like a serial tty driver with a KISS TNC attached. This means that configuration of the AX.25 software for this driver is identical to the configuration for a KISS TNC presented earlier. You need to use the axattach program and add the port to your /etc/ax25/axports file. The only difference is that you would use the device /dev/bc0 instead of /dev/ttyS0.

14 Configuring ax25d to accept incoming AX.25/Netrom connections.

Linux is a powerful operating system and offers a great deal of flexibility in how it is configured. With this flexibility comes a cost in configuring it to do what you want. When configuring your Linux machine to accept incoming AX.25 or NetRom connections there are a number of questions you need to ask yourself. The most important of which is: ”What do I want users to see when they connect?”. As detailed in the HAM-HOWTO there are BBS systems that are being developed that you might want users to see when they connect, alternatively you might want to give users a login prompt so that they can make use of a shell account, or you might even have written your own program, such as a customised database or a game, that you want people to connect to. Whatever you choose, you must tell the AX.25 software about this so that it knows what software to run when it accepts an incoming AX.25 connection.

14.1 Creating the /etc/ax25/ax25d.conf file.

This file is the configuration file for the ax25d AX.25 daemon which handles incoming AX.25 and NetRom connections.

The file is a little cryptic looking at first, but you’ll soon discover it is very simple in practice, with a small trap for you to be wary of.

The format of the ax25d.conf file is as follows:

     # This is a comment and is ignored by the ax25d program.
     [<interface_call>]
     <peer1>  window T1 T2 T3 N2 <mode> <uid> <cmd> <cmd-name> <arguments>
     <peer2>  window T1 T2 T3 N2 <mode> <uid> <cmd> <cmd-name> <arguments>
     defaults window T1 T2 T3 N2 <mode> <uid> <cmd> <cmd-name> <arguments>
     <peer3>  window T1 T2 T3 N2 <mode> <uid> <cmd> <cmd-name> <arguments>
     ...
     default  window T1 T2 T3 N2 <mode> <uid> <cmd> <cmd-name> <arguments>

Where:

# at the start of a line marks a comment and is completely ignored by the ax25d program.

<interface call> is the AX.25 callsign of the AX.25 or NetRom interface that this particular paragraph is for.

<peer> is the callsign of the peer node that this particular configuration applies to. If you don’t specify an SSID here then any SSID will match.

window is the AX.25 Window parameter (K) or MAXFRAME parameter for this configuration.

T1

is the Frame retransmission (T1) timer in half second units. T2

is the amount of time the AX.25 software will wait for another incoming frame before preparing a response in 1 second units.

T3 is the amount of time of inactivity before the AX.25 software will disconnect the session in 1 second units.

N2

is the number of consecutive retransmissions that will occur before the connection is closed. <mode>

provides a mechanism for determining certain types of general permissions. The modes are enabled or disabled by supplying a combination of characters, each representing a permission. The characters may be in either upper or lower case and must be in a single block with no spaces. u/U

UTMP - currently unsupported. v/V

Valid call - currently unsupported. q/Q

Quiet - Don’t log connection n/N

check NetRom Neighbour - currently unsupported. d/D

Disallow Digipeaters - Connections must be direct, not digipeated. l/L

Lockout - Don’t allow connection. */0

marker - place marker, no mode set.

<uid>

is the userid that the program to be run to support the connection should be run as. <cmd>

is the full pathname of the command to be run, with no arguments specified. <cmd-name>

is the text that should appear in a ps as the command name running (normally the same as <cmd> except without the directory path information.

<arguments> are the command line argument to be passed to the <:cmd> when it is run. You pass useful information into these arguments by use of the following tokens:

%U

AX.25 callsign of the connected party without the SSID, in uppercase. %u

AX.25 callsign of the connected party without the SSID, in lowercase. %S

AX.25 callsign of the connected party with the SSID, in uppercase. %s

AX.25 callsign of the connected party with the SSID, in lowercase.

You need one section in the above format for each AX.25 or NetRom interface you want to accept incoming AX.25 or NetRom connections on.

There are two special lines in the paragraph, one starts with the string ‘defaults’ and other starts with the string ‘default’ (yes there is a difference). These lines serve special functions.

The ‘default’ lines purpose should be obvious, this line acts as a catch-all, so that any incoming connection on the <interface call> interface that doesn’t have a specific rule will match the ‘default’ rule. If you don’t have a ‘default’ rule, then any connections not matching any specific rule will be disconnected immediately without notice.

The ‘defaults’ line is a little more subtle, and here is the trap I mentioned earlier. In any of the fields for any definition for a peer you can use the ‘*’ character to say ‘use the default value’. The ‘default’ line is what sets those default values. The kernel software itself has some defaults which will be used if you don’t specify any. The trap is that the these defaults apply only to those rules below the ‘default’ line not to those above. You may have more than one ‘default’ rule per interface definition, and in this way you may create groups of default configurations.

14.2 A simple example ax25d.conf file.

Ok, an illustrative example:

     # ax25d.conf for VK2KTJ - 21/11/95
     # This configuration uses the AX.25 port defined earlier.
     [VK2KTJ-0]
     NOCALL     *     *  *  *  *   L    guest /usr/games/fortune
     #
     defaults   1    10  *  *  *   *    root  /usr/local/sbin/axspawn axspawn %u +
     VK2XLZ-1   *     *  *  *  *   *      *     *         *
     VK2DAY-1   *     *  *  *  *   *      *     *         *
     default    1    10  5 100 5   *    root  /usr/local/bin/pms pms -a -o vk2ktj
     #

This example says that anybody connecting to the interface on my machine with the callsign VK2KTJ-0 will have the following rules applied:

Anyone whose callsign is set to ‘NOCALL’ should use the kernel default parameters and have the fortune program run as guest for them.

The defaults line changes two parameters from the kernel defaults (Window and T1) and will run the /usr/local/sbin/axspawn program for them. Any copies of /usr/local/sbin/axspawn run this way will appear as axspawn in a ps listing for convenience. The next two lines definitions for two stations who will receive those permissions.

The last line in the paragraph is the ‘catch all’ definition that everybody else will get (including VK2XLZ and VK2DAY using any other SSID other than -1). This definition sets all of the parameters implicitly and will cause the pms program to be run with a command line argument indicating that it is being run for an AX.25 connection (assuming VK2KTJ-0 to be an AX.25 interface) and that the owner callsign is VK2KTJ. (See the ‘Configuring the PMS’ section below for more details).

If you wanted to define rules for another interface, a NetRom interface for example, then you would add another paragraph coded in the same way with rules and callsign appropriate to it.

This example is a contrived one but I think it illustrates clearly the important features of the syntax of the configuration file. Jonathon includes a much longer and more detailed example in the ax25-utils package that you should look at if you need more information.

14.3 Starting ax25d

When you have the two configuration files completed you start ax25d with the command:

     # /usr/local/sbin/ax25d

When this is run people should be able to make AX.25 connections to your Linux machine. Remember to put the ax25d command in your rc files so that it is started automatically when you reboot each time.

14.4 Configuring the node software.

The node software was developed by Tomi Manninen <tomi.manninen@hut.fi> and was based on the original PMS program. It provides a fairly complete and flexible node capability that is easily configured. It allows users once they are connected to make telnet, netrom or telnet connections out and to obtain various sorts of information such as Finger, Nodes and Heard lists etc.

The node would normally be invoked from the ax25d program although it is also capable of being invoked from the tcp/ip inetd program to allow users to telnet to your machine and obtain access to it, or by running it from the command line.

14. Configuring ax25d to accept incoming AX.25/Netrom connections. 33

14.4.1 Creating the /usr/local/node/node.conf file.

The node.conf file is where the main configuration of the node takes place. It is a simple text file and its format is as follows:

# /usr/local/node/node.conf
# configuration file for the node(8) program.
#
# Lines beginning with '#' are comments and are ignored.
# Hostname
# Specifies the hostname of the node machine
hostname        radio.gw.vk2ktj.ampr.org
# Local Network
# allows you to specify what is consider 'local' for the
# purposes of permission checking using nodes.perms.
localnet        44.136.8.96/29
# Hide Ports
# If specified allows you to make ports invisible to users. The
# listed ports will not be listed by the (P)orts command.
hiddenports     70cm 2m
# Callserver
# If configured the node provides users with access to a callserver.
callserver      zone.oh7rba.ampr.org
# Node Idle Timeout
# Specifies the idle time for connections to this node in seconds.
idletimout      1800
# Connection Idle Timeout
# Specifies the idle timer for connections made via this node in
# seconds.
conntimeout     1800
# Reconnect
# Specifies whether users should be reconnected to the node
# when their remote connections disconnect, or whether they
# should be disconnected complete.
reconnect       on
# Command Aliases
# Provide a way of making complex commands simple.
alias           CONV    "telnet vk1xwt.ampr.org 3600"
alias           BBS     "connect radio vk2xsb"

14.4.2 Creating the /usr/local/node/node.perms file.

The node allows you to assign permissions to users. These permissions allow you to determine which users should be allowed to make use of options such as the (T)elnet, and (C)onnect commands, for example, and which shouldn’t. The node.perms file is where this information is stored and contains five key fields. For all fields an asterisk ‘*’ character matches anything. This is useful for building default rules.

user The first field is the callsign or user to which the permissions should apply. Any SSID value is ignored, so you should just place the base callsign here.

method Each protocol or access method is also given permissions. For example you might allow users who have connected via AX.25 or NetRom to use the (C)onnect option, but prevent others, such as those who are telnet connected from a non-local node from having access to it. The second field therefore allows you to select which access method this permissions rule should apply to. The access methods allowed are:

     method  description
     ------  -----------------------------------------------------------
     ax25    User connected by AX.25
     netrom  User connected by NetRom
     host    User started node from command line
     local   User is telnet connected from a 'local' host
     ampr    User is telnet connected from an amprnet address (44.0.0.0)
     inet    user is telnet connected from a non-loca, non-ampr address.

port For AX.25 users you can control permissions on a port by port basis too if you choose. This allows you to determine what AX.25 are allowed to do based on which of your ports they have connected to. The third field contains the port name if you are using this facility. This is useful only for AX.25 connections.

password You may optionally configure the node so that it prompts users to enter a password when they connect. This might be useful to help protect specially configured users who have high authority levels. If the fourth field is set then its value will be the password that will be accepted.

permissions The permissions field is the final field in each entry in the file. The permissions field is coded as a bit field, with each facility having a bit value which if set allows the option to be used and if not set prevents the facility being used. The list of controllable facilities and their corresponding bit values are:

     value   description
     -----   -------------------------------------------------

           1      Login allowed.
           2      AX25 (C)onnects allowed.
           4      NetRom (C)onnects allowed.
           8      (T)elnet to local hosts allowed.
           16     (T)elnet to amprnet (44.0.0.0) hosts allowed.
           32     (T)elnet to non-local, non-amprnet hosts allowed.
           64     Hidden ports allowed for AX.25 (C)onnects.

To code the permissions value for a rule, simply take each of the permissions you want that user to have and add their values together. The resulting number is what you place in field five.

A sample nodes.perms might look like:

     # /usr/local/node/node.perms
     #
     # The node operator is VK2KTJ, has a password of 'secret' and
     # is allowed all permissions by all connection methods
     vk2ktj  *       *       secret  127
     # The following users are banned from connecting
     NOCALL  *       *       *       0
     PK232   *       *       *       0
     PMS     *       *       *       0
     # INET users are banned from connecting.
     *       inet    *       *       0
     # AX.25, NetRom, Local, Host and AMPR users may (C)onnect and (T)elnet
     # to local and ampr hosts but not to other IP addresses.
     *       ax25    *       *       31
     *       netrom  *       *       31
     *       local   *       *       31
     *       host    *       *       31
     *       ampr    *       *       31

14.4.3 Configuring node to run from ax25d

The node program would normally be run by the ax25d program. To do this you need to add appropriate rules to the /etc/ax25/ax25d.conf file. In my configuration I wanted users to have a choice of either connecting to the node or connecting to other services. ax25d allows you to do this by cleverly creating creating port aliases. For example, given the ax25d configuration presented above, I want to configure node so that all users who connect to VK2KTJ-1 are given the node. To do this I add the following to my

/etc/ax25/ax25d.conf file:
     [VK2KTJ-1 VIA VK2KTJ-0]
     default         * *    *   *   *   0    root /usr/local/bin/node node radio

This simply says that all users who connect to VK2KTJ-1 on the same device (by the same port) as VK2KTJ-0 should cause the node program to be run with the name of the port they have connected via, in this case you will recall the /etc/ax25/axports specifies the port name as radio that supports the callsign VK2KTJ-0.

14.4.4 Configuring node to run from inetd

If you want users to be able to telnet a port on your machine and obtain access to the node you can go this fairly easily. The first thing to decide is what port users should connect to. In this example I’ve arbitrarily chosen port 4000, though Tomi gives details on how you could replace the normal telnet daemon with the node in his documentation.

You need to modify two files.

To /etc/services you should add:

     node    4000/tcp        #OH2BNS's node software

and to /etc/inetd.conf you should add:

     node    stream  tcp     nowait  root    /usr/local/bin/node node

When this is done, and you have restarted the inetd program any user who telnet connects to port 4000 of your machine will be prompted to login and if configured, their password and then they will be connected to the node.

14.5 Configuring axspawn.

The axspawn program is a simple program that allows AX.25 stations who connect to be log in to your machine. It is invoked from the ax25d program as described above. To allow a user to log in to your machine you should add a line similar to the following into your /etc/ax25/ax25d.conf file:

     default * * * * * 1 root /usr/local/sbin/axspawn axspawn %u

If the line ends in the + character then the connecting user must hit return before they will be allowed to login. The default is to not wait. Any individual host configurations that follow this line will have the axspawn program run when they connect. When axspawn is run it first checks that the command line argument it is supplied is a legal callsign, strips the SSID, then it checks that /etc/passwd file to see if that user has an account configured. If there is an account, and the password is either "" (null) or + then the user is logged in, if there is anything in the password field the user is prompted to enter a password. If there is not an existing account in the /etc/passwd file then axspawn may be configured to automatically create one.

14.5.1 Creating the /etc/ax25/axspawn.conf file.

You can alter the behaviour of axspawn in various ways by use of the /etc/ax25/axspawn.conf file. This file is formatted as follows:

     # /etc/ax25/axspawn.conf
     #
     # allow automatic creation of user accounts
     create    yes
     #
     # guest user if above is 'no' or everything else fails. Disable with "no"
     guest     no
     #
     # group id or name for autoaccount
     group     ax25
     #
     # first user id to use
     first_uid 2001
     #
     # maximum user id
     max_uid   3000
     #
     # where to add the home directory for the new users
     home      /home/ax25
     #
     # user shell
     shell     /bin/bash
     #
     # bind user id to callsign for outgoing connects.
     associate yes

The eight configurable characteristics of axspawn are as follows:

#

indicates a comment.

create

if this field is set to yes then axspawn will attempt to automatically create a user account for any user who connects and does not already have an entry in the /etc/passwd file.

guest

this field names the login name of the account that will be used for people who connect who do not already have accounts if create is set to no. This is usually ax25 or guest.

group

this field names the group name that will be used for any users who connect and do not already have an entry in the /etc/passwd file. first uid

this is the number of the first userid that will be automatically created for new users.

max uid this is the maximum number that will be used for the userid of new users.

home this is the home (login) directory of new users.

shell this is the login shell of any new users.

associate this flag indicates whether outgoing AX.25 connections made by this user after they login will use their own callsign, or your stations callsign.

14.6 Configuring the pms

The pms program is an implementation of a simple personal message system. It was originally written by Alan Cox. Dave Brown, N2RJT, <dcb@vectorbd.com> has taken on further development of it. At present it is still very simple, supporting only the ability to send mail to the owner of the system and to obtain some limited system information but Dave is working to expand its capability to make it more useful.

Configuring the pms program is still a little clumsy at the moment because it is new but bare with it while Dave works on it. The only difficult part is modifying the source code to support your local mail delivery agent and whether you use the ttylinkd program or the talk-ax25 program.

After that is done there are a couple of simple files that you should create that give users some information about the system and then you need to add appropriate entries into the ax25d.conf file so that connected users are presented with the PMS.

14.6.1 Customising pms.c

If you take a look at the pms.c file in the ax25-utils package you will see that there a few items that you can configure. Perhaps the most important to configure is the one that allows you to specify what local mail delivery agent you use. In most installations the deliver program may well work, but in the minimal installation on my gateway machine I needed to modify the line that previously said:

#define MAIL_DELIVERY_AGENT     "deliver %s"

to:

#define MAIL_DELIVERY_AGENT     "/bin/mail.local %s"

There is no easy way of knowing what local delivery agent your machine is configured to use, so perhaps you could simply look for either of these files in the obvious places and choose whichever seems to be there.

I also have the ttylinkd program installed, so I needed to modify the section relating to it. I keep my copy of the ttylinkd program in my /usr/sbin directory, so I modified the line that previously said:

/* #define MY_TALK              "/usr/local/sbin/ttylinkd" */

to:

#define MY_TALK              "/usr/sbin/ttylinkd"

If you use the talk-ax25 program then you would remove the /* and */ from around that line instead.

There is one last definition that might be useful to you, and that is if you have a default email gateway. This line is the line containing the string MY GATEWAY. If you have an email gateway that you use, then you should remove the comments from this line and change the address to that of your gateway.

14.6.2 Create the /usr/local/pms/motd file.

The /usr/local/pms/motd file contains the ‘message of the day’ that users will be presented with after they connect and receive the usual BBS id header. The file is a simple text file, any text you include in this file will be sent to users.

14.6.3 Create the /usr/local/pms/info file.

The /usr/local/pms/info file is also a simple text file in which you would put more detailed information about your station or configuration. This file is presented to users in response to their issuing of the Info command from the PMS> prompt.

14.6.4 Associate AX.25 callsigns with system users.

When a connected user sends mail to an AX.25 callsign, the pms expects that callsign to be mapped, or associated with a real system user on your machine.

You make this association with the axparms command. An example looks like:

     # axparms -assoc vk2ktj terry

This command associates that AX.25 callsign vk2ktj with the user terry on the machine, so any mail for vk2ktj on the pms will be sent to terry.

Remember to put these associations into your rc file so that they are available each time your reboot.

14.6.5 Add the PMS to the /etc/ax25/ax25d.conf file.

Adding the pms to your ax25d.conf file is very simple. There is one small thing you need to think about though. Dave has added command line arguments to the PMS to allow it to handle a number of different text end-of-line conventions. AX.25 and NetRom by convention expect the end-of-line to be carriage return, linefeed while the standard unix end-of-line is just newline. So, for example, if you wanted to add an entry that meant that the default action for a connection received on an AX.25 port is to start the PMS then you would add a line that looked something like:

     default  1  10 5 100 5   0    root  /usr/local/bin/pms pms -a -o vk2ktj

This simply runs the pms program, telling it that it is an AX.25 connection it is connected to and that the PMS owner is vk2ktj. Check the man page for what you should specify for other connection methods.

14.6.6 Test the PMS.

To test the PMS, you can try the following command from the command line:

# /usr/local/bin/pms -u vk2ktj -o vk2ktj

Substitute your own callsign for mine and this will run the pms, telling it that it is to use the unix end-of-line convention, and that user logging in is vk2ktj. You can do all the things connected users can.

Additionally you might try getting some other node to connect to you to confirm that your ax25d.conf configuration works.

15 HOWTO link NOS and the Linux kernel networking software

Many people like to run some version of NOS under Linux because it has all of the features and facilities they are used to. Most of those people would also like to have the NOS running on their machine capable of talking to the Linux kernel so that they can offer some of the linux capabilities to radio users.

15.1 Linking NOS and Linux using a pipe device

Brandon S. Allbery, KF8NH, contributed the following information to explain how to interconnect the NOS running on a Linux machine with the kernel code using the Linux pipe device.

Since both Linux and NOS support the slip protocol it is possible to link the two together by creating a slip link. You could do this by using two serial ports with a loopback cable between them, but this would be slow and costly. Linux provides a feature that many other Unix-like operating systems provide called ‘pipes’. These are special pseudo devices that look like a standard tty device to software but in fact loopback to another pipe device. To use these pipes the first program must open the master end of the pipe, and the open then the second program can open the slave end of the pipe. When both ends are open the programs can communicate with each other simply by writing characters to the pipes in the way they would if they were terminal devices.

To use this feature to connect the Linux Kernel and a copy of NOS, or some other program you first must choose a pipe device to use. You can find one by looking in your /dev directory. The master end of the pipes are named: ptyp[1-f] and the slave end of the pipes are known as: ttyp[1-f]. Remember they come in pairs, so if you select /dev/ptypf as your master end then you must use /dev/ttypf as the slave end.

Once you have chosen a pipe device pair to use you should allocate the master end to you linux kernel and the slave end to the NOS program, as the Linux kernel starts first and the master end of the pipe must be opened first. You must also remember that your Linux kernel must have a different IP address to your NOS, so you will need to allocate a unique address for it if you haven’t already.

You configure the pipe just as if it were a serial device, so to create the slip link from your linux kernel you can use commands similar to the following:

     # /sbin/slattach -s 38400 -p slip /dev/ptypf &
     # /sbin/ifconfig sl0 broadcast 44.255.255.255 pointopoint 44.70.248.67 /
             mtu 1536 44.70.4.88
     # /sbin/route add 44.70.248.67 sl0
     # /sbin/route add -net 44.0.0.0 gw 44.70.248.67

In this example the Linux kernel has been given IP address 44.70.4.88 and the NOS program is using IP address 44.70.248.67. The route command in the last line simply tells your linux kernel to route all datagrams for the amprnet via the slip link created by the slattach command. Normally you would put these commands into your /etc/rc.d/rc.inet2 file after all your other network configuration is complete so that the slip link is created automatically when you reboot. Note: there is no advantage in using cslip instead of slip as it actually reduces performance because the link is only a virtual one and occurs fast enough that having to compress the headers first takes longer than transmitting the uncompressed datagram.

To configure the NOS end of the link you could try the following:

     # you can call the interface anything you want; I use "linux" for convenience.
     attach asy ttypf - slip linux 1024 1024 38400
     route addprivate 44.70.4.88 linux

These commands will create a slip port named ‘linux’ via the slave end of the pipe device pair to your linux kernel, and a route to it to make it work. When you have started NOS you should be able to ping and telnet to your NOS from your Linux machine and vice versa. If not, double check that you have made no mistakes especially that you have the addresses configured properly and have the pipe devices around the right way.

16 The /proc/ file system entries.

The /proc filesystem contains a number of files specifically related to the AX25 and NetRom kernel software. These files are normally used by the AX52 utilities, but they are plainly formatted so you may be interested in reading them. The format is fairly easily understood so I don’t think much explanation will be necessary.

arp contains the list of Address Resolution Protocol mappings of IP addresses to MAC layer protocol addresses. These can can AX.25, ethernet or some other MAC layer protocol.

ax25

contains the AX.25 port configuration information. ax25%lowbar;bpqether

contains the AX25 over ethernet BPQ style callsign mappings. ax25 calls

contains the linux userid to callsign mappings set my the axparms -assoc command. ax25 route

contains AX.25 digipeater path information. nr

contains the NetRom port configuration information. nr neigh

contains information about the NetRom neighbours known to the NetRom software. nr nodes

contains information about the NetRom nodes known to the NetRom software.

17 Walking on the wilder side .. the developmental software

If all of the above doesn’t satisfy your desire to push the limits there is plenty more for you to play with. There have a been a number of enhancements that are available as experimental software for you to try. They come either as patches against the 2.0.* kernels, or in the alpha 2.1.* kernels. Be warned though, this software is new, pretty much untested and likely to contain bugs. If that worries you, then ignore this section of the document.

17.1 Enhanced versions of just about everything.

The 2.1.* kernels have enhanced versions of nearly all of the protocols and drivers. The most significant of the enhancements are:

modularised the protocols and drivers have all been modularised so that you can insmod and rmmod them whenever you wish. This reduces the kernel memory requirements for infrequently used modules and makes development and bug hunting much simpler.

all drivers are now network drivers all of the network devices such as Baycom, SCC, PI, Packettwin etc now present a normal network interface, that is they now look like the ethernet driver does, they no longer look like KISS TNC’s. A new utility called net2kiss allows you to build a kiss interface to these devices if you wish.

bug fixed there have been many bug fixes and new features added to the drivers and protocols.

To take advantage of the features of the 2.1.* AX.25 support you must use the new ax25-utils package. You shouldn’t use a kernel older than 2.1.9 and you should use ax25-utils-2.1.9a.tar.gz which is available from Jonathons home site.

17.2 ROSE Packet Layer Support

The Rose packet layer protocol is similar to layer three of the X.25 specification. The kernel based Rose support is a modified version of the FPAC Rose implementation

<http://fpac.lmi.ecp.fr/f1oat/f1oat.html>. You will need ax25-utils-2.1.9a.tar.gz to configure

the rose interface.

17.3 Soundcard DSP based modem

Thomas Sailor has built a new driver for the kernel that allows you to use your soundcard as a modem. Connect your radio directly to your soundcard to play packet!

18 Some sample configurations.

Following are examples of the most common types of configurations. These are guides only as there are as many ways of configuring your network as there are networks to configure, but they may give you a start.

18.1 Small Ethernet LAN with Linux as a router to Radio LAN

18. Some sample configurations. 44

Many of you may have small local area networks at home and want to connect the machines on that network to your local radio LAN. This is the type of configuration I use at home. I arranged to have a suitable block of addresses allocated to me that I could capture in a single route for convenience and I use these on my Ethernet LAN. Your local IP coordinator will assist you in doing this if you want to try it as well. The addresses for the Ethernet LAN form a subset of the radio LAN addresses. The following configuration is the actual one for my linux router on my network at home:

                                          .      .   .    .    . .
  ---                                .
   | Network       /---------\     .    Network
   | 44.136.8.96/29|         |    .     44.136.8/24        \ | /
   |               | Linux   |   .                          \|/
   |               |         |  .                            |
   |          eth0 | Router  |  .  /-----\    /----------\   |
   |---------------|         |-----| TNC |----| Radio    |---/
   |   44.136.8.97 |  and    |  .  \-----/    \----------/
   |               |         | sl0
   |               | Server  | 44.136.8.5
   |               |         |    .
   |               |         |     .
   |               \_________/       .
  ---                                     .      .   .    .    . .
#!/bin/sh
# /etc/rc.net
# This configuration provides one KISS based AX.25 port and one
# Ethernet device.
echo "/etc/rc.net"
echo "  Configuring:"
echo -n "    loopback:"
/sbin/ifconfig lo 127.0.0.1
/sbin/route add 127.0.0.1
echo " done."
echo -n "    ethernet:"
/sbin/ifconfig eth0 44.136.8.97 netmask 255.255.255.248 \
                broadcast 44.136.8.103 up
/sbin/route add 44.136.8.97 eth0
/sbin/route add -net 44.136.8.96 netmask 255.255.255.248 eth0
echo " done."
echo -n "    AX.25: "
axattach -i 44.136.8.5 -m 512 /dev/ttyS1 4800
ifconfig sl0 netmask 255.255.255.0 broadcast 44.136.8.255
route add -host 44.136.8.5 sl0
route add -net 44.136.8.0 window 1024 sl0
echo -n "    Netrom: "

  nrattach -i 44.136.8.5 netrom
  echo "  Routing:"
  /sbin/route add default gw 44.136.8.68 window 1024 sl0
  echo "    default route."
  echo done.
  # end
/etc/ax25/axports
  # name  callsign        speed   paclen  window  description
  4800    VK2KTJ-0        4800    256     2       144.800 MHz
/etc/ax25/nrports
  # name  callsign        alias   paclen  description
  netrom  VK2KTJ-9        LINUX   235     Linux Switch Port
/etc/ax25/nrbroadcast
  # ax25_name     min_obs def_qual        worst_qual      verbose
  4800            1       120             10              1

• You must have IP FORWARDING enabled in your kernel.

• The AX.25 configuration files are pretty much those used as examples in the earlier sections, refer to those where necessary.

• I’ve chosen to use an IP address for my radio port that is not within my home network block. I needn’t have done so, I could have easily used 44.136.8.97 for that port too.

• 44.136.8.68 is my local IPIP encapsulated gateway and hence is where I point my default route.

• Each of the machines on my Ethernet network have a route:

       route add -net 44.0.0.0 gw 44.136.8.97 window 512 mss 512 eth0

The use of the mss and window parameters means that I can get optimum performance from both local Ethernet and radio based connections.

• I also run my sendmail, http, ftp and other daemons on the router machine so that it needs to be the only machine to provide others with facilities.

• The router machine is a lowly 386DX20 with a 20Mb harddrive and a very minimal linux configuration.

18.2 IPIP encapsulated gateway configuration.

Linux is now very commonly used for tcp/ip encapsulated gateways around the world. The new tunnel driver supports multiple encapsulated routes and makes the older ipip daemon obsolete.

A typical configuration would look similar to the following.

                                          .      .   .    .    . .
  ---                                .
   | Network       /---------\     .    Network
   | 154.27.3/24   |         |    .     44.136.16/24       \ | /
   |               | Linux   |   .                          \|/
   |               |         |  .                            |
   |          eth0 | IPIP    |  .  /-----\    /----------\   |
---|---------------|         |-----| TNC |----| Radio    |---/
   |   154.27.3.20 | Gateway |  .  \-----/    \----------/
   |               |         | sl0
   |               |         | 44.136.16.1
   |               |         |    .
   |               |         |     .
   |               \_________/       .
  ---                                     .      .   .    .    . .

The configuration files of interest are:

# /etc/rc.net
# This file is a simple configuration that provides one KISS AX.25
# radio port, one Ethernet device, and utilises the kernel tunnel driver
# to perform the IPIP encapsulation/decapsulation
#
echo "/etc/rc.net"
echo "  Configuring:"
#
echo -n "    loopback:"
/sbin/ifconfig lo 127.0.0.1
/sbin/route add 127.0.0.1
echo " done."
#
echo -n "    ethernet:"
/sbin/ifconfig eth0 154.27.3.20 netmask 255.255.255.0 \
                broadcast 154.27.3.255 up
/sbin/route add 154.27.3.20 eth0
/sbin/route add -net 154.27.3.0 netmask 255.255.255.0 eth0
echo " done."
#
echo -n "    AX.25: "
axattach -i 44.136.16.1 -m 512 /dev/ttyS1 4800
/sbin/ifconfig sl0 netmask 255.255.255.0 broadcast 44.136.16.255

/sbin/route add -host 44.136.16.1 sl0
/sbin/route add -net 44.136.16.0 netmask 255.255.255.0 window 1024 sl0
#
echo -n "    tunnel:"
/sbin/ifconfig tunl0 44.136.16.1 mtu 512 up
#
echo done.
#
echo -n "Routing ... "
source /etc/ipip.routes
echo done.
#
# end.

and:

# /etc/ipip.routes
# This file is generated using the munge script
#
/sbin/route add -net 44.134.8.0 netmask 255.255.255.0 tunl0 gw 134.43.26.1
/sbin/route add -net 44.34.9.0 netmask 255.255.255.0 tunl0 gw 174.84.6.17
/sbin/route add -net 44.13.28.0 netmask 255.255.255.0 tunl0 gw 212.37.126.3
   ...
   ...
   ...
/etc/ax25/axports
# name  callsign        speed   paclen  window  description
4800    VK2KTJ-0        4800    256     2       144.800 MHz

Some points to note here are:

• The new tunnel driver uses the gw field in the routing table in place of the pointopoint parameter to specify the address of the remote IPIP gateway. This is why it now supports multiple routes per interface.

• You can configure two network devices with the same address. In this example both the sl0 and the tunl0 devices have been configured with the IP address of the radio port. This is done so that the remote gateway sees the correct address from your gateway in encapsulated datagrams sent to it.

• The route commands used to specify the encapsulated routes can be automatically generated by a modified version of the munge script. This is included below. The route commands would then be written to a separate file and read in using the bash source /etc/ipip.routes command (assuming you called the file with the routing commands /etc/ipip.routes) as illustrated. The source file must be in the NOS route command format.

• Note the use of the window argument on the route command. Setting this parameter to an appropriate value improves the performance of your radio link.

The new tunnel-munge script:

#!/bin/sh
#
# From: Ron Atkinson <n8fow@hamgate.cc.wayne.edu>
#
#  This script is basically the 'munge' script written by Bdale N3EUA
#  for the IPIP daemon and is modified by Ron Atkinson N8FOW. It's
#  purpose is to convert a KA9Q NOS format gateways route file
#  (usually called 'encap.txt') into a Linux routing table format
#  for the IP tunnel driver.
#
#        Usage: Gateway file on stdin, Linux route format file on stdout.
#               eg.  tunnel-munge < encap.txt > ampr-routes
#
# NOTE: Before you use this script be sure to check or change the
#       following items:
#
#     1) Change the 'Local routes' and 'Misc user routes' sections
#        to routes that apply to your own area (remove mine please!)
#     2) On the fgrep line be sure to change the IP address to YOUR
#        gateway Internet address. Failure to do so will cause serious
#        routing loops.
#     3) The default interface name is 'tunl0'. Make sure this is
#        correct for your system.
echo "#"
echo "# IP tunnel route table built by $LOGNAME on `date`"
echo "# by tunnel-munge script v960307."
echo "#"
echo "# Local routes"
echo "route add -net 44.xxx.xxx.xxx netmask 255.mmm.mmm.mmm dev sl0"
echo "#"
echo "# Misc user routes"
echo "#"
echo "# remote routes"
fgrep encap | grep "^route" | grep -v " XXX.XXX.XXX.XXX" | \
awk '{
        split($3, s, "/")
        split(s[1], n,".")
        if      (n[1] == "")    n[1]="0"
        if      (n[2] == "")    n[2]="0"
        if      (n[3] == "")    n[3]="0"
        if      (n[4] == "")    n[4]="0"

if      (s[2] == "1")   mask="128.0.0.0"
else if (s[2] == "2")   mask="192.0.0.0"
else if (s[2] == "3")   mask="224.0.0.0"
else if (s[2] == "4")   mask="240.0.0.0"
else if (s[2] == "5")   mask="248.0.0.0"
else if (s[2] == "6")   mask="252.0.0.0"
else if (s[2] == "7")   mask="254.0.0.0"
else if (s[2] == "8")   mask="255.0.0.0"
else if (s[2] == "9")   mask="255.128.0.0"
else if (s[2] == "10")  mask="255.192.0.0"
else if (s[2] == "11")  mask="255.224.0.0"
else if (s[2] == "12")  mask="255.240.0.0"
else if (s[2] == "13")  mask="255.248.0.0"
else if (s[2] == "14")  mask="255.252.0.0"
else if (s[2] == "15")  mask="255.254.0.0"
else if (s[2] == "16")  mask="255.255.0.0"
else if (s[2] == "17")  mask="255.255.128.0"
else if (s[2] == "18")  mask="255.255.192.0"
else if (s[2] == "19")  mask="255.255.224.0"
else if (s[2] == "20")  mask="255.255.240.0"
else if (s[2] == "21")  mask="255.255.248.0"
else if (s[2] == "22")  mask="255.255.252.0"
else if (s[2] == "23")  mask="255.255.254.0"
else if (s[2] == "24")  mask="255.255.255.0"
else if (s[2] == "25")  mask="255.255.255.128"
else if (s[2] == "26")  mask="255.255.255.192"
else if (s[2] == "27")  mask="255.255.255.224"
else if (s[2] == "28")  mask="255.255.255.240"
else if (s[2] == "29")  mask="255.255.255.248"
else if (s[2] == "30")  mask="255.255.255.252"
else if (s[2] == "31")  mask="255.255.255.254"
else                    mask="255.255.255.255"
if (mask == "255.255.255.255")
printf "route add -host %s.%s.%s.%s gw %s dev tunl0\n"\
        ,n[1],n[2],n[3],n[4],$5
else
printf "route add -net %s.%s.%s.%s gw %s netmask %s dev tunl0\n"\
        ,n[1],n[2],n[3],n[4],$5,mask
}'
echo "#"
echo "# default the rest of amprnet via mirrorshades.ucsd.edu"
echo "route add -net 44.0.0.0 gw 128.54.16.18 netmask 255.0.0.0 dev tunl0"
echo "#"
echo "# the end"

19 Discussion relating to Amateur Radio and Linux.

There are various places that discussion relating to Amateur Radio and Linux take place. They take place in the comp.os.linux.* newsgroups, they also take place on the HAMS list on vger.rutgers.edu. Other places where they are held include the tcp-group mailing list at ucsd.edu (the home of amateur radio tcp/ip discussions), and you might also try the #linpeople channel on the linuxnet irc network.

To join the Linux linux-hams channel on the mail list server, send mail to:

     Majordomo@vger.rutgers.edu

with the line:

     subscribe linux-hams

in the message body. The subject line is ignored.

To join the tcp-group send mail to:

     listserver@ucsd.edu

with the line:

     subscribe tcp-group

in the body of the text.

Note: Please remember that the tcp-group is primarily for discussion of the use of advanced protocols, of which tcp/ip is one, in Amateur Radio. Linux specific questions should not ordinarily go there.

20 Copyright.

The AX25-HOWTO, information on how to install and configure some of the more important packages providing AX25 support for Linux. Copyright (c) 1996 Terry Dawson.

This program is free software; you can redistribute it and/or modify it under the terms of the GNU General Public License as published by the Free Software Foundation; either version 2 of the License, or (at your option) any later version.

This program is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABILITY or FITNESS FOR A PARTICULAR PURPOSE. See the GNU General Public License for more details.

You should have received a copy of the GNU General Public License along with this program; if not, write to the:

Free Software Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.